Form an equation whose roots are the reciprocals of the roots of x^4-5x^3+7x^2-3x+6=0
1
Expert's answer
2021-08-10T13:21:04-0400

Put "\frac{1}{x}" in the given equation



"\frac{1}{x^4}-\frac{5}{x^3}+\frac{7}{x^2}-\frac{3}{x}+6=0"



"\frac{1-5x+7x^2-3x^3+6x^4}{x^4}=0"



"6x^4-3x^3+7x^2-5x+1=0"